Kayressa and John both drive delivery trucks. They leave the warehouse at the same time, but their routes are in opposite directions. Kayressa's rate is 15 mph less than John's rate. After 4 hours, they are 380 miles apart. What are the rates of the trucks?
Kayressa's rate = ____ mph
and
John's rate = _____ mph.

Answer by rfer(12662)

380/4=95
x+x-15=95
2x=110
x=55 mph john
x-5=40 mph kay